OpenStudy (anonymous):

What is the solution to the equation –13 + x = –4? A. –9 B. 9 C. 17 D. –17

OpenStudy (anonymous):

Have you ever had to work with equations like this?

OpenStudy (anonymous):

Add 13 to both sides to solve it

OpenStudy (anonymous):

the solution is B -13+x=-4 x=-4+13 x=9